# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

### v2

@@
Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities, but still struggle with issues like hallucinations and outdated information. Retrieval-augmented generation (RAG) addresses these issues by grounding LLM outputs in external knowledge with an Information Retrieval (IR) system. Building on this foundation, graph-based RAG systems go a step further by retrieving subgraphs, which preserve the relationships between knowledge entities and provide more comprehensive context. However, graph RAG faces two challenges: (1) Retrieving relevant information introduces irrelevant nodes (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and leads to overly lengthy inputs that hinder efficiency; (2) The representation gap between graph and language during generation with LLMs limits the ability to fully leverage graph structures for enhanced understanding. To address these limitations, we propose Align-GRAG, a novel reasoning-guided dual alignment framework in post-retrieval phrase. It first formulates a subgraph by retrieving nodes and edges. Then an Aligner is proposed to jointly optimize a graph encoder with an LLM-summarized reasoning chain. It achieves dual alignment of graph node and representation by leveraging KL divergence loss and contrastive loss, facilitating efficient pruning of irrelevant knowledge and establishing a unified semantic space. The Generator integrates the aligned graph data with LLM to produce coherent and accurate answers. Experiments on the GraphQA benchmark across three tasks (including common sense reasoning, scene graph understanding, and knowledge graph reasoning) validate the effectiveness of our method. The codes are available in this repository [^1] https://anonymous.4open.science/r/Align-GRAG-F3D8/.
Retrieval-Augmented Generation, Large Language Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
@@
Recent advancements in natural language processing have demonstrated the remarkable capabilities of large language models (LLMs), such as GPT (OpenAI, 2024) and Llama (Touvron et al., 2023), in understanding, reasoning, and handling complex tasks (Zhao et al., 2023; Xu et al., 2024a). The massive parameterized knowledge encoded within LLMs has enabled them to excel in information retrieval (Zhu et al., 2023; AlKhamissi et al., 2022; Petroni et al., 2019). Despite these advancements, concerns about the interpretability of LLM remain, leading to challenges like hallucinations (fabrication of false or misleading information) (Huang et al., 2023; Xu et al., 2024c) and reliance on outdated data (Gao et al., 2023). These issues are especially concerning in high-stakes fields such as healthcare (He et al., 2023) and law (Lai et al., 2024).
Retrieval-augmented generation (RAG) systems have been developed to solve these problems (Gao et al., 2023). By integrating external information retrieval (IR) systems, RAG retrieves knowledge from external databases, ensuring that LLMs are equipped with up-to-date and relevant information tailored to the user’s query. By grounding the LLM’s outputs in verifiable, external knowledge, RAG systems significantly improve the reliability and accuracy of content generated by LLMs (Fan et al., 2024). Nevertheless, in real-world scenarios, RAG systems often divide long context into independent chunks, overlooking the deeper connections between fragments and lacking a global perspective (Edge et al., 2024; Guo et al., 2024). Moreover, a significant portion of data inherently exhibits a graph-like structure, such as in recommendation systems (Guo et al., 2020), the Web, and knowledge graphs (Pan et al., 2024). To address these, researchers have proposed graph-based RAG (GRAG) (Edge et al., 2024; He et al., 2024; Tian et al., 2024; Sun et al., [n. d.]) as an enhancement to traditional RAG methods. Instead of retrieving isolated text chunks where information is treated independently of its structural relationships, GRAG retrieves subgraphs from databases (He et al., 2024). These subgraphs maintain the interconnected relationships between knowledge entities, providing richer context and enabling more accurate, contextually relevant responses.
Figure 1. Comparison of Align-GRAG with traditional RAG (including post-retrieval methods such as reranker). Left: query and textual graph (entities and relations). Middle (a): Traditional RAG retrieves top‑k passages and may use a reranker, but still suffers from noisy/irrelevant evidence and loses the underlying graph structure, which hurts answer grounding. Right (b): Align‑GRAG introduces a graph aligner that jointly performs node alignment and representation alignment, preserving structural relations and filtering irrelevant content, showcasing its effectiveness in graph RAG scenarios.
However, despite its potential, integrating RAG with graphs is non-trivial, presenting two key challenges: (1) Irrelevant Knowledge: The retrieval stage often introduces irrelevant information, hindering the LLM’s attention on useful knowledge (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and the generated inputs become overly lengthy, further causing computational inefficiency. As shown in Figure 1(a), some traditional RAG approaches (Xiao et al., 2024; Li et al., 2023b; Jia et al., 2024) aim to tackle this challenge by incorporating a reranker in the post-retrieval stage (Fan et al., 2024; Gao et al., 2023). For instance, BGE (Xiao et al., 2024) leverages cross-encoder-based architectures to effectively capture interactions between queries and documents, enabling more precise relevance estimation. However, in the GRAG setting, where chunks are typically represented as nodes or triplets, this approach can disrupt structural information and overlook the connections between knowledge. Meanwhile, existing GRAG methods (Guo et al., 2024) focus primarily on retrieval-stage optimizations (e.g., GraphRAG (Edge et al., 2024) builds community graphs for long documents to facilitate retrieval and summarization), resulting in relatively shallow similarity that lacks deeper semantic matching capabilities. (2) Representation Gap: In the generation stage, when integrating LLMs with graph structure embeddings, such as those generated by Graph Neural Network (GNN), a challenge arises due to the representation gap: graph embeddings are designed to capture structural properties, which misalign with the sequence-based representations used by LLMs. This misalignment limits the ability to fully leverage the graph for enhanced reasoning (Liu et al., 2023; Zhao et al., 2022). In the field of LLMs, most existing GRAG methods (Sun et al., [n. d.]; LUO et al., [n. d.]; Xu et al., 2025; Hu et al., 2022; Chen et al., 2024; Jiang et al., 2024) primarily treat graph data as plain text inputs, lacking a dedicated encoder for structural information, which limits their understanding capabilities. Some methods (Tian et al., 2024; Perozzi et al., 2024; He et al., 2024) incorporate graph encoders, but typically rely on straightforward approaches like concatenating graph embeddings with LLM inputs. For instance, G-Retriever (He et al., 2024) uses a projector to map graph embeddings, while GNP (Tian et al., 2024) applies cross-modality pooling to fuse text and graph representations. However, neither approach explicitly aligns these distinct representation spaces, leaving the underlying gap unresolved. These research gaps present a question: Can we better prune irrelevant nodes while keeping structural information, and effectively encode graph representations for utilization by LLMs?
In this work, we propose Align-GRAG, a novel reasoning-guided dual alignment framework specifically designed for GRAG. During the retrieval stage, we calculate the similarity between nodes and edges to extract the initial subgraph. To address the two key challenges in GRAG, we introduce an innovative graph aligner module for the post-retrieval phase, as shown in Figure 1(b). We leverage LLM-based summarization ability with well-crafted prompts to generate reasoning chains that bridge information gaps. The aligner achieves dual alignment with reasoning chains as optimization target: (1) Graph Node Alignment, which ranks critical reasoning nodes, pruning irrelevant knowledge; and (2) Graph Representation Alignment, ensuring unified semantic space between graph and language representations via contrastive learning. This dual alignment refines subgraphs to focus on relevant knowledge, enabling the generator to produce accurate, context-aware responses grounded in the graph’s structure. Extensive experiments are conducted on GraphQA benchmark (He et al., 2024), covering tasks such as commonsense reasoning, scene graph understanding, and knowledge graph reasoning, which demonstrate the superior performance of Align-GRAG compared to other baselines and highlight the effectiveness of Aligner module. Overall, this work makes three key contributions:
  • To the best of our knowledge, this is the first work to introduce a reasoning-guided graph alignment framework in the post-retrieval phrase for GRAG.
  • We propose a novel graph aligner module for dual alignment between graph node and representation, allowing effective pruning of irrelevant nodes while learning a unified semantic space.
  • The effectiveness of Align-GRAG was comprehensively validated on three datasets across different tasks, demonstrating superior performance compared to other strong baselines.
2. Related Work
2.1. Retrieval-Augmented Generation (RAG)
RAG (Fan et al., 2024; Gao et al., 2023; Fan et al., 2024) has been extensively studied to address the challenges of hallucination and outdated information (Huang et al., 2023; Xu et al., 2024c), which often lead to untrustworthy outputs from LLMs (Lewis et al., 2020). By integrating information retrieval systems, RAG has demonstrated its effectiveness in real-world applications, enhancing reliability of generated responses (Salemi and Zamani, 2024; Zamani and Bendersky, 2024; Yang et al., 2024; Sudhi et al., 2024). The RAG pipeline mainly includes: pre-retrieval (e.g., query rewriting and query expansion), retrieval, post-retrieval (e.g., reranking), and generation (Fan et al., 2024). Among these, reranking has emerged as a key technique for refining retrieval results. Initial retrieval (Robertson et al., 2009; Izacard et al., 2021) often relies on simple similarity scoring. Reranking (Li et al., 2023b; Xiao et al., 2024; Dong et al., 2024), on the other hand, employs more sophisticated models to reassess and reorder the initially retrieved documents based on their relevance to the query, thereby improving the quality and relevance of retrieved documents. For instance, bge-reranker (Xiao et al., 2024) leverages cross-encoders that perform full attention over the input pair, providing higher accuracy compared to embedding-based models. However, in GRAG scenarios, reranking may disrupt the inherent structural information of graph. Our proposed Align-GRAG effectively performs post-retrieval processing to extract relevant knowledge, while preserves the graph structural information.
2.2. Large Language Models on Graph
Graphs, composed of nodes and edges, are essential for modeling real-world relationships across various domains. In recent years, GNN (Shi et al., 2020; Veličković et al., 2018; Kipf and Welling, 2016; Wu et al., 2020) have emerged as a powerful tool for encoding graph structures. With the rise of LLMs and their demonstrated impressive capabilities, there is growing interest in integrating LLMs with graph learning techniques to enhance graph-based tasks (Huang et al., 2024b; Jin et al., 2024; Li et al., 2023a; Ren et al., 2024). Pioneering works (Tang et al., 2024; Perozzi et al., 2024; Chai et al., 2023) have incorporated GNNs by feeding graph tokens into LLMs. Beyond basic integration, recent studies (Qin et al., 2023; Zhu et al., 2024; Huang et al., 2024a) have focused on tighter architectural fusion, embedding graph neural layers within transformer architectures to enable seamless interaction between graph reasoning and natural language understanding . However, most existing researches (Ren et al., 2024) focus on tasks like node and graph classification. Our work addresses the challenges of retrieval-based graph QA, leveraging retrieval-augmented methods and graph learning techniques while overcoming the limitations of current frameworks.
2.3. Graph RAG
Traditional RAG often struggles to capture structured relational knowledge or global context. Graph Retrieval-Augmented Generation (GRAG) addresses these limitations by leveraging graph-structured knowledge for improved retrieval and reasoning (Baek et al., 2023; Wang et al., 2024; Han et al., 2024; Xu et al., 2024b; Zhang et al., 2025). GRAG utilizes graph databases (e.g., Freebase (Bollacker et al., 2008), Wikidata (Vrandečić and Krötzsch, 2014)) to retrieve graph elements like triples, paths, or subgraphs (Edge et al., 2024; Mavromatis and Karypis, 2024; Hu et al., 2024; Wu et al., 2024; Xiang et al., 2025; Xiao et al., 2025). Some approaches (Edge et al., 2024; Guo et al., 2024; Fan et al., 2025; Liang et al., 2024) focus on constructing large-scale graphs from text. For example, GraphRAG (Edge et al., 2024) uses LLMs to extract entities, partitions graphs using hierarchical clustering, and generates summaries to support retrieval. Other methods (Yasunaga et al., 2021; Jiang et al., 2022; He et al., 2024; Tian et al., 2024) encode graph data with GNNs and integrate it with language models for enhanced reasoning, while some translate natural language queries into logical forms for knowledge graph retrieval (Zhang et al., 2023; Yu et al., 2022; Luo et al., 2023; Xu et al., 2025; Hu et al., 2022). More recently, LLMs have been used as iterative agents for reasoning over knowledge graphs (Jiang et al., 2024; LUO et al., [n. d.]; Sun et al., [n. d.]; Chen et al., 2024; Zhang et al., 2024). However, existing GRAG approaches primarily emphasize optimizing retrieval but often lack effective post-retrieval strategies, leading to shallow subgraph-query similarity. To address this, our Align-GRAG introduces a novel dual alignment mechanism that bridges retriever and generator, improving overall performance.
3. Preliminaries
This section presents relevant concepts and notations, including the definition of textual graphs and the task formalization.
Textual Graph. A textual graph refers to a graph where the nodes and edges are supplemented with textual information, enabling the representation of both structural relationships and semantic content. Formally, a textual graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E},\{t_{n}\}_{n\in\mathcal{V}},\{t_{e}\}_{e\in\mathcal{E}})$, where, $\mathcal{V}$ and $\mathcal{E}$ represent the sets of nodes and edges, respectively. $t_{n}\in\mathcal{D}^{L_{n}}$ denotes the sequential text associated with a node $n\in\mathcal{V}$, where $\mathcal{D}$ is the vocabulary set, and $L_{n}$ is the length of the text associated with the node. Similarly, $t_{e}\in\mathcal{D}^{L_{e}}$ represents the sequential text associated with an edge $e\in\mathcal{E}$, and $L_{e}$ denotes the length of the text associated with the edge.
Task Formulation. This work focuses on RAG with textual graph databases. The goal is to retrieve relevant information from the structured nature of textual graphs and generate coherent, contextually accurate responses. Formally, a given query $t_{q}=[x_{1},...,x_{n}]$, represented as a sequence of tokens from the vocabulary $\mathcal{D}$. The model retrieves relevant subgraphs $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$ from $\mathcal{G}$ by computing the semantic similarity between the query $t_{q}$ and the node and edge texts $t_{n}$ and $t_{e}$. where $\mathcal{V}_{r}\subseteq\mathcal{V}$ and $\mathcal{E}_{r}\subseteq\mathcal{E}$. The retrieval process identifies relevant nodes, edges, and their associated texts in the context of the input question $t_{q}$. In the post-retrieval stage, we focus on refining the retrieved information to enhance the quality of input provided to LLM, ensuring more accurate and contextually relevant outputs. In the generation stage, the retrieved subgraph, along with the query $t_{q}$ and prompt $\mathcal{P}$, is provided as input to LLM. The LLM generates the final output $Y=[y_{1},...,y_{m}]$, grounded in the retrieved knowledge and the contextual understanding of the graph structure.
4. Methodology
In this section, we introduce the Align-GRAG framework, which is composed of three components: Graph Retriever (Sec. 4.1), Graph Aligner (Sec. 4.2), and Graph Generator (Sec. 4.3) modules. The overall framework pipeline is depicted in Figure 2. In Figure 1, ➀ denotes Graph Retriever, we leverage the similarity between entities, relations, and the query to extract the initial subgraph. In the aligner module (Figure 2-➁), we use a GNN backbone to encode subgraph information. To tackle irrelevant knowledge noise and representation gaps, we introduce an LLM-based summarization technique with tailored prompts. By taking the query-answer pair and subgraph as input, the LLM generates a reasoning chain that links the query to the answer, identifying key intermediate nodes and bridging information gaps. Next, we propose the graph aligner to achieve dual alignment with this summarized reasoning chain: (1) Graph Node Alignment: A KL divergence loss (Kullback and Leibler, 1951) is utilized to highlight critical nodes and reasoning edges, aligning node-level distributions and pruning irrelevant nodes. (2) Graph Representation Alignment: A contrastive loss with in-batch negative sampling is further designed to align graph and language representations in a shared semantic space. Jointly optimizing these alignments enables effective knowledge pruning and unified representation learning. In the generation stage (Figure 2-➁), the aligned graph embedding is used as a graph token, concatenated with the query and pruned subgraph, and then fed into an LLM for answer generation. By leveraging the aligned graph, generator produces responses that reflect a deep understanding of graph’s structure and its relevance to query.
Figure 2. The Align-GRAG framework includes three components: ➀ Graph Retriever, ➁ Graph Aligner, and ➂ Graph Generator. It first retrieves subgraphs, then trains the aligner with an LLM-based summarized reasoning chain. The trained aligner prunes irrelevant nodes, generates graph embeddings, and feeds them into the LLM to produce the final answer.
4.1. Graph Retriever
Existing RAG methodologies are mainly designed for plain text documents or triplets, where information retrieval occurs independently of the graph structure (Fan et al., 2024; Zhu et al., 2023; Pan et al., 2024). In the retrieval stage, we first utilize an encoder-only language model (e.g., SentenceBERT (Reimers and Gurevych, 2020)) to encode textual information, including the query $t_{q}$, text of each node $t_{n}$ and edge $t_{e}$ in the graph, respectively:
(1) $\boldsymbol{q}$ $=\text{SBERT}(t_{q})\in\mathbb{R}^{d},$
(2) $\boldsymbol{n}$ $=\text{SBERT}(t_{n})\in\mathbb{R}^{d},$
(3) $\boldsymbol{e}$ $=\text{SBERT}(t_{e})\in\mathbb{R}^{d}$
(4) $\mathcal{V}_{k}$ $=\operatorname*{arg\,topk}_{\begin{subarray}{c}\boldsymbol{n}\in\mathcal{V}\end{subarray}}\big(\text{sim}(\boldsymbol{q},\boldsymbol{n})\big),$
(5) $\mathcal{E}_{k}$ $=\operatorname*{arg\,topk}_{\begin{subarray}{c}\boldsymbol{e}\in\mathcal{E}\end{subarray}}\big(\text{sim}(\boldsymbol{q},\boldsymbol{e})\big)$
This forms top-$k$ entities and relation sets, denoted as $\mathcal{G}_{\text{retriever}}$. Inspired by He et al. (2024), we further leverage the Prize-Collecting Steiner Tree algorithm (PCST) (Bienstock et al., 1993) to maintain a controlled graph size.
@@ 4.1. Graph Retriever
(7) $\mathcal{G}_{\text{retriever}}=\underset{\begin{subarray}{c}S\subseteq\mathcal{G},\\ S\text{ is connected}\end{subarray}}{\arg\max}\left(\sum_{n\in\mathcal{V}_{S}}\text{prize}(n)+\sum_{e\in\mathcal{E}_{S}}\text{prize}(e)-\text{cost}(S)\right)$
where $\mathcal{V}_{S}$ and $\mathcal{E}_{S}$ are the sets of nodes and edges in the subgraph $S$, respectively. The cost of constructing the subgraph is defined as $\text{cost}(S)=|\mathcal{E}|\cdot C_{e}$, where $|\mathcal{E}|$ is the number of edges, and $C_{e}$ is a predefined per-edge cost that serves as a regularization parameter to control the subgraph’s size. In this way, we can obtain a preliminary retrieved subgraph $\mathcal{G}_{\text{Retriever}}$.
4.2. Graph Aligner
In this section, we introduce the Graph Aligner module. The retrieved subgraph $\mathcal{G}_{\text{Retriever}}$ faces limitations when integrated with LLMs, including irrelevant information from the retriever and misalignment between graph and language representations during generation with LLM. Our Graph Aligner tackles these issues through two key objectives: (1) Aligning node-level distributions to prune irrelevant structures, refining subgraphs based on LLM knowledge preferences; (2) Bridging the gap between graph structures and language descriptions by unifying them in a shared latent space for LLM integration. The illustration is shown in Figure 2-➁. We provide a detailed explanation below.
4.2.1. Summarized Reasoning Chain
To achieve the Aligner module’s optimization goals, our motivation is to leverage a crucial reasoning chain related to the query-answer pair. This is especially important for multi-hop problems requiring intermediate reasoning steps to bridge information gaps. Incorporating such reasoning significantly improves answer quality, as demonstrated by methods like Chain-of-Thoughts (Wei et al., 2022) and O1 (Zhong et al., 2024). To explicitly identify the essential nodes within graph database, we propose an LLM-based summarization technique. By designing well-crafted prompts, we harness the capabilities of strong LLMs to generate comprehensive reasoning chains that logically connect the question to the answer. For example, Table 6 in the case study, given a query, “What is the name of the first Harry Potter novel?”, with the LLM-generated prior, we can identify the intermediate and critical node, J.K. Rowling. We believe this can serve as an excellent label. The deeper insights provided through reasoning are subsequently utilized in the dual node and representation alignment process to train the aligner module. More details for the summarization prompt are provided in experiments 5.10.
4.2.2. Node Alignment

### v3

# Align-GRAG: Anchor and Rationale Guided Dual Alignment for Graph Retrieval-Augmented Generation

@@
Abstract
Despite the strong abilities, large language models (LLMs) still suffer from hallucinations and reliance on outdated knowledge, raising concerns in knowledge-intensive tasks. Graph-based retrieval-augmented generation (GRAG) enriches LLMs with knowledge by retrieving graphs leveraging relational evidence, but it faces two challenges: structure-coupled irrelevant knowledge introduced by neighbor expansion and structure-reasoning discrepancy between graph embeddings and LLM semantics. We propose Align-GRAG, an anchor-and-rationale guided refinement framework to address these challenges. It prompts an LLM to extract anchors and rationale chains, which provide intermediate supervision for (1) node-level alignment that identifies critical nodes and prunes noisy evidence, and (2) graph-level alignment that bridges graph and language semantic spaces via contrastive learning. Extensive experiments on commonsense reasoning, scene graph understanding, and knowledge graph reasoning demonstrate consistent gains over 18 strong baselines, validating the effectiveness of Align-GRAG for improving graph-grounded generation. The code can be found in https://anonymous.4open.science/r/Align-GRAG-F3D8/.
Align-GRAG: Anchor and Rationale Guided Dual Alignment for Graph Retrieval-Augmented Generation
Derong Xu*1,2, Pengyue Jia*2, Xiaopeng Li2, Yingyi Zhang2, Maolin Wang2, Qidong Liu2, Xiangyu Zhao2, Yichao Wang3, Huifeng Guo3, Ruiming Tang3, Enhong Chen1, Tong Xu1 1University of Science and Technology of China, 2City University of Hong Kong, 3Noah’s Ark Lab, Huawei *Equal contribution.
1 Introduction
@@
Recent progress in large language models (LLMs) has highlighted their strong abilities in understanding, reasoning Zhao et al. (2023); Xu et al. (2024a), and knowledge-intensive tasks Zhu et al. (2023); AlKhamissi et al. (2022); Sui et al. (2025). Despite these advances, LLMs still raise reliability concerns, including hallucinations (producing false or misleading information) Huang et al. (2023); Xu et al. (2024c) and reliance on outdated data Fang et al. (2025). These issues are especially concerning in high-stakes fields such as healthcare He et al. (2023) and law Lai et al. (2024).
Retrieval-augmented generation (RAG) addresses these concerns by retrieving query-relevant evidence from external databases and grounding generation in a verifiable context, improving factuality and robustness Fan et al. (2024); Zhou et al. (2025). Nevertheless, in real-world scenarios, RAG systems often divide long context into independent chunks, overlooking the deeper connections between fragments and lacking a global perspective Guo et al. (2024). Moreover, many knowledge sources are inherently graph-structured, including recommendation systems Wang et al. (2025), the Web, and knowledge graphs Pan et al. (2024). Motivated by this, graph-based RAG (GRAG) Edge et al. (2024); He et al. (2024); Tian et al. (2024); Sun et al. extends RAG by retrieving subgraphs rather than isolated passages He et al. (2024), preserving correlations among multiple text units.
Figure 1: The challenges in graph RAG scenarios.
However, integrating RAG with graphs is non-trivial, raising two GRAG-specific challenges, as shown in Figure 1. (I) Structure-Coupled Irrelevant Knowledge. To ensure multi-hop coverage, GRAG typically does expansions with retrieved interconnected neighbors. While necessary, this process introduces “structure-coupled” noise–irrelevant nodes that are topologically close to the target but semantically unhelpful for the specific reasoning chain. Crucially, determining the utility of these nodes requires a holistic view of the reasoning path rather than assessing local query similarity. Existing post-retrieval rerankers Xiao et al. (2024); Li et al. (2023); Jia et al. (2024), designed for independent text segments, fail in this context: by treating nodes or triples as isolated units, they sever the structural dependencies essential for judging reasoning-dependent relevance. As a result, current GRAG systems Guo et al. (2024); Edge et al. (2024) lack a structure-aware post-retrieval refinement mechanism with sufficient semantic modeling. (II) Structure-Reasoning Discrepancy. Incorporating structure embeddings (e.g., from GNNs) is hindered by a representation gap: graph embeddings primarily capture structural signals, whereas LLMs operate over sequence semantics. This mismatch makes it difficult for LLM reasoning to reliably exploit graph structure, limiting graph-grounded generation Liu et al. (2023); Zhao et al. (2022). Existing GRAG methods either linearize graphs as text Sun et al. ; LUO et al. ; Xu et al. (2025); Hu et al. (2022); Chen et al. (2024); Jiang et al. (2024) or incorporate graph encoders with simple concatenation/projection Tian et al. (2024); Perozzi et al. (2024); He et al. (2024); however, they typically lack an explicit objective that aligns structural representations with the semantics required for LLM reasoning. These research gaps present a question:
[Uncaptioned image]How can we construct a compact yet structure-preserving subgraph and learn aligned graph representations for better LLM reasoning?
@@
Figure 2: Overview of Align-GRAG. We first extract a rationale chain and anchors with LLM. The rationale guides graph-level alignment, while anchors supervise node-level alignment.
2 Preliminaries
Textual Graph.
A textual graph is a graph where nodes and edges are enriched with textual information, capturing structural and semantic details. A textual graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E},\{t_{n}\}_{n\in\mathcal{V}},\{t_{e}\}_{e\in\mathcal{E}})$, where $\mathcal{V}$ and $\mathcal{E}$ are the sets of nodes and edges. $t_{n}\in\mathcal{D}^{L_{n}}$ represents the text associated with a node $n\in\mathcal{V}$, where $\mathcal{D}$ is the vocabulary, and $L_{n}$ is the length. Similarly, $t_{e}\in\mathcal{D}^{L_{e}}$ is the text for edge $e\in\mathcal{E}$ with length $L_{e}$.
Task Formulation.
This work addresses RAG with textual graph. The goal is to retrieve relevant information from textual graphs and generate accurate responses. Given a query $t_{q}$, a sequence of tokens from $\mathcal{D}$, the model retrieves a subgraph $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$ from $\mathcal{G}$ based on the semantic similarity of $t_{n}$, $t_{e}$ with $t_{q}$, where $\mathcal{V}_{r}\subseteq\mathcal{V}$ and $\mathcal{E}_{r}\subseteq\mathcal{E}$. The retrieved nodes, edges, and texts are refined to improve the input quality for the LLM. During generation, the retrieved subgraph, query $t_{q}$, and prompt $\mathcal{P}$ are provided as input to the LLM, which produces the final output $Y$, grounded in the retrieved knowledge and graph context.
3 Methodology
This section presents Align-GRAG (Figure 2), an anchor-and-rationale guided refinement framework for GRAG. Align-GRAG first extracts anchors concepts and compact rationale chains with LLM, and then uses them to guide a graph aligner that (i) performs structure-aware pruning to remove irrelevant evidence while preserving relational dependencies and (ii) reduces the gap between graph and language representations.
3.1 Anchor and Rationale Extraction
To optimize the Aligner module, we leverage an explicit reasoning signal extracted for each query, which is especially important for multi-hop questions requiring intermediate steps. Inspired by Chain-of-Thought Wei et al. (2022) and O1-style reasoning Zhong et al. (2024), we propose an LLM-based extraction module that produces (i) a compact rationale chain and (ii) a small set of anchors.
Specifically, given a question $q$ and its answer, we first retrieve an initial subgraph (as elaborated in Appendix A). We then prompt a strong LLM to extract (i) a concise rationale chain that connects $q$ to the supporting evidence and (ii) anchors, i.e., key intermediate entities/relations that can be grounded to graph nodes, from the retrieved subgraph. For example, in Figure 9, for the query “What is the name of the first Harry Potter novel?”, the extracted rationale highlights J.K. Rowling as a critical intermediate concept, which serves as an anchor. These outputs support dual alignment: the anchors node-level alignment for structure-aware pruning, while rationale guides supervise graph-level alignment to better ground language representations to graph.
3.2 Guided Dual Alignment
Node-level Alignment.